# PAX-TS: Model-agnostic multi-granular explanations for time series forecasting via localized perturbations

### v1

@@ I Introduction
Time series forecasting models have seen a wave of research over the last years, with subsequent performance improvements, where new generations of models outperform older models. Recently, this has mainly been driven by transformer architectures [1, 2, 3] and foundation models [4], promising high performance across different forecasting tasks. Forecasting models have also been applied in the industry [^1] https://www.uber.com/blog/forecasting-introduction/, providing business value for many companies. However, when models are applied to real-time tasks, practitioners and domain experts commonly face challenges in comprehending how and why a model arrived at its forecast. To resolve this issue, explainable artificial intelligence (XAI) [5] is a paradigm that is becoming highly relevant for time series forecasting tasks. However, this problem still remains underexplored, with existing work falling short of providing appropriate and meaningful explanations for AI-based forecasts in time series applications.
@@
State-of-the-art time series forecasting models ([6, 7, 2, 8, 3]) are generally opaque, providing no explanations for their forecasts, making them less intuitive for the end-user. To combat this, existing post-hoc XAI methods like LIME [9], which were developed for classification tasks, have been applied to forecasting [10, 11]. However, explanation methods designed for classification are generally not suitable for this task because a forecasting model predicts multiple values over time, in contrast to classification models, which output a single class label for each data point.
@@
In this work, we propose Perturbation Analysis eXplanations for Time Series forecasting (PAX-TS), a technique for multi-granular human-comprehensible explanations, specific to time series forecasting. PAX-TS uses locally perturbed inputs to assess how a forecasting model generates its forecasts, giving insights at different levels of granularity, up to the time-step-level. Fig. 1 shows an example of input perturbations ($\mathbf{x}^{\prime}$) of the global minimum of an individual sample ($\mathbf{x}$) and the respective forecasts ($\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime}$) for each perturbed input with different scale parameters (+5% and -10%).
@@
    Novelty: We introduce PAX-TS, a post-hoc model-agnostic XAI technique designed for time series forecasting. PAX-TS can be used on both univariate and multivariate time series, and its explanations can be generated at different granularities, ranging from high-level cross-channel correlations to low-level time-step-importance explanations.
@@
    Empirical evaluation: We conduct a diverse benchmark with 7 time series forecasting algorithms on 10 datasets. We further compare the explanations of PAX-TS to the explanations of TS-MULE and ShapTime.
@@
    Performance explanation: Based on the benchmark results and the explanations obtained by PAX-TS, we derive a set of visually distinct temporal dependency patterns that are an indicator of algorithmic performance.
  4. 4.
    Complexity analysis: We assess the time complexity of PAX-TS, which can be used in real-time to provide explanations for practitioners working with the data.
  5. 5.
    Reproducibility: We describe PAX-TS and its algorithm in detail and provide reproducible code in the following publicly accessible repository:
    https://anonymous.4open.science/r/pax-ts-6410
@@ II Background
Explainability for time series forecasts can be achieved in two ways: Inherently explainable models are algorithms that are designed to be more comprehensible, providing both the forecast and additional information, which explains the forecast in different ways. Secondly, post-hoc methods can be applied based on the trained model and its forecast to explain how the model has generated the respective forecast. Baur et al. [12] provide a meta-review on explainability for electric load forecasting, summarising a range of techniques used in this application area. In this section, we describe both inherently explainable forecasting algorithms and post-hoc explanation techniques to give a broad overview of the related work.
Algorithms providing inherent explanations. N-BEATS [13] is a state-of-the-art univariate time series forecasting algorithm based on multilayer perceptron stacks, which additionally offers a more interpretable architecture. In this architecture, the provided explanations are based on trend and seasonality, which sum up to the model’s forecast. This simple yet effective approach allows the end-user to get a more fine-grained understanding of the model’s forecast by dividing the forecast into components, which are simpler to comprehend.
In their work, Wang et al. [14] propose a wavelet decomposition-based LSTM forecaster named mLSTM, which provides explanations based on decomposed components. The authors demonstrate feature importance explanations for each time step of the input window for each component resulting from the multilevel discrete wavelet transform. The importance scores are calculated based on relative input changes, which show how much the overall model prediction changes. However, this is irrespective of whether the change in model prediction is beneficial or not.
A small cluster of work utilizes fuzzy logic for forecasting tasks [15, 16, 17]. In the review of Baur et al. [12], explanations resulting from fuzzy logic membership functions are visualized. When applied on a larger scale, fuzzy rules lead to cognitive maps, linking features of the dataset to each other. However, the related work does not delve deeper into the explainability provided by fuzzy logic for time series forecasting, and this topic remains largely unexplored.
@@
Lim et al. [1] introduce temporal fusion transformers, an attention-based forecasting algorithm utilizing past data, static covariates, and known future inputs to make multivariate forecasts. Their algorithm provides values of relative importance to each of the input features. This results in an overall comparison, demonstrating which feature affects the forecast the most. However, this only explains which feature affects a forecast, whether in a positive or negative way, only in a specific segment, or for certain peaks.
In a similar way, Pantiskas et al. [18] use the attention mechanism to generate feature importance scores by time step for their proposed forecasting model. With this method, a specific output time step can be related to the entire input window, showing the importance of each input time step for this specific time step. This forecast explanation method is cumbersome, especially with multivariate data and longer input and output window lengths.
An alternative solution to provide inherently explainable time series forecasts is probabilistic forecasting, as used in the work of Roth et al. [19]. The authors use a state space model to quantify posterior uncertainty, which allows for confidence intervals for each time step of the prediction. This helps illustrate the model’s confidence for the forecast and allows the end-user to more clearly grasp the algorithm’s accuracy.
Post-hoc explanation techniques. Zhang et al. [11] propose ShapTime, a method that adapts Shapley Values [20] to characterize the importance of segments in the input sequence. With this method, the importance of each input segment on the mean of the forecast can be assessed, effectively describing how much a given segment affects the forecast overall. While this is useful for some use cases, the granularity of the explanation is too coarse to provide actionable insights.
Troncoso-García et al. [21] propose a method for discovering association rules for univariate time series forecasting. The resulting association rules link single values in the input and the output window with intervals of varying size. This type of explanation is problematic, as many association rules are found, and, in addition to that, their intervals are often large, providing little practical value for the end-user.
@@
TS-MULE [10] is a post-hoc explanation technique for the time series forecasting task, utilizing the perturbation-based approach of LIME [9] to make explanations. TS-MULE includes different segmentation techniques to generate explanations, resulting in feature importance scores for segments of the input window. This approach faces problems similar to other feature importance-based techniques like SHAP or attention maps, as the importance of a feature relates to the accuracy of the entire forecast, which is not sufficient to explain time series forecasts.
@@ III The PAX-TS algorithm
For the multivariate time series forecasting problem, a $d$-dimensional time series subsequence of length $b$, $\mathbf{x}\in\mathbb{R}^{d\times b}$, is to be forecast for a horizon of length $h$, resulting in the forecast $\hat{\mathbf{y}}\in\mathbb{R}^{d\times h}$, aiming to approximate the ground truth $\mathbf{y}\in\mathbb{R}^{d\times h}$. Note that $\mathbf{x}$ and $\mathbf{y}$ are consecutive, and the problem is referred to as univariate time series forecasting when $d=1$. In the literature, $b$ and $h$ vary, however, we have designed our method with a focus on short-term forecasting with $h\leq 48$, to effectively visualize explanations. Given an input subsequence $\mathbf{x}$, a trained forecasting model $M_{\theta}(\cdot)$ generates a forecast $\hat{\mathbf{y}}$, with more accurate forecasting models yielding a closer approximation to the ground truth $\mathbf{y}$, resulting in lower forecasting error.
Our proposed approach to explaining time series forecasts, PAX-TS, is perturbation-based and focuses on the relation between input and output subsequence on an index level based on summary statistics and indices of interest. In sections III-A-III-C, we outline how realistic perturbations can be applied to indices and summary statistics. Section III-D describes how to apply PAX-TS and calculate change ratios, section III-E analyzes its complexity, and section III-F details PAX-TS’ usage for cross-channel correlations of multivariate data.
@@ III-A Localized perturbation by index
A localized Gaussian-smoothed perturbation is applied to perturb a positive time series subsequence $\mathbf{x}$ of length $b$ at a given time step $t$, denoted as $\mathbf{x}_{t}$, by a scale parameter $\alpha$, where $\alpha>0$ indicates a shift in upward direction, and $\alpha<0$ indicates a shift in downward direction. We introduce this smoothing operation to achieve more plausible perturbations compared to point-based perturbations. Additionally, a width parameter $w\in\mathbb{N}$ affects the perturbation’s range, and a softness parameter $s\in\mathbb{R}$ influences the Gaussian fall-off.
To apply the perturbation, we first calculate a positional weight $w_{p}$ for Gaussian-like smoothing, calculated for each time step $i$ of the subsequence. The weight is based on the distance between time step $i$ and $t$, and it is set to 0 for all time steps outside the fixed window size $w$:
@@
$w_{a}\;=\;\frac{\min(\mathbf{x}_{i},\mathbf{x}_{t})}{\max(\mathbf{x}_{i},\mathbf{x}_{t})+\epsilon},\qquad\forall\,i\in\{1,\dots,b\}$
@@
$\mathbf{x}^{\prime}\;=\;\mathbf{x}_{i}\;+\;w_{p}w_{a}\alpha\mathbf{x}_{i},\qquad\forall\,i\in\{1,\dots,b\}$
@@ III-B Summary statistic scaling
In this work, we consider the first and second moments of the time series as relevant summary statistics for scaling due to their simplicity and practical relevance. The same procedure can be followed for higher-order moments, however, the equations need to be adapted accordingly.
With a scale parameter $\alpha$, the first moment of a time series subsequence $\mathbf{x}$ can be scaled as follows:
$\mathbf{x^{\prime}\;=\;\mathbf{x}+\alpha\mu},\qquad\mu=\frac{1}{b}\sum_{i=1}^{b}\mathbf{x_{i}}$
@@
$\mathbf{x}^{\prime}\;=\;(\mathbf{x}-\mu)*\sqrt{\alpha+1}+\mu,\qquad\mu=\frac{1}{b}\sum_{i=1}^{b}\mathbf{x_{i}}$
@@ III-C Trend-drift adjustment
The drift, representing the overall trend of a time series subsequence $\mathbf{x}$, can be adjusted by deseasonalizing the sequence, fitting a linear polynomial, and adjusting its slope according to a scale parameter $\alpha$.
Given a seasonality length $k\in\mathbb{N}$, a discrete convolution operation, denoted by ”*”, results in a deseasonalized moving average of the time series:
@@
Using ordinary least squares, a linear polynomial can be fit to the residual $\mathbf{x}_{d}$, resulting in an intercept $a$ and a slope $m$, representing the trend of the deseasonalized sequence. The adjusted slope can be obtained as follows:
@@
With the adjusted slope, the intercept can be modified in three different ways, resulting in different adjustments:
Right-fixed: $a^{\prime}\;=\;a-(b-1)z$
Symmetric: $a^{\prime}\;=\;a-\dfrac{1}{2}(b-1)z$
Left-fixed: $a^{\prime}\;=\;a$

### v2

@@ I Introduction
Time series forecasting models have seen a wave of research over the last years, with subsequent performance improvements, where new generations of models outperform older models. Recently, this has mainly been driven by transformer architectures [lim2021temporal, liu2023itransformer, naghashi2025multiscale] and foundation models [das2024decoder], promising high performance across different forecasting tasks. Forecasting models have also been applied in the industry [^1] https://www.uber.com/blog/forecasting-introduction/, providing business value for many companies. However, when models are applied to real-time tasks, practitioners and domain experts commonly face challenges in comprehending how and why a model arrived at its forecast. To resolve this issue, explainable artificial intelligence (XAI) [angelov2021explainable] is a paradigm that is becoming highly relevant for time series forecasting tasks. However, this problem still remains underexplored, with existing work falling short of providing appropriate and meaningful explanations for AI-based forecasts in time series applications.
@@
State-of-the-art time series forecasting models [lin2023segrnn, liu2023itransformer, wang2024timemixer, naghashi2025multiscale] are generally opaque, providing no explanations for their forecasts, making them less intuitive for the end-user. To combat this, existing post-hoc XAI methods like LIME [ribeiro2016should], which were developed for classification tasks, have been applied to forecasting [schlegel2021ts, zhang2023shaptime]. However, explanation methods designed for classification are generally not suitable for this task because a forecasting model predicts multiple values over time, in contrast to classification models, which output a single class label for each data point.
@@
In this work, we propose Perturbation Analysis eXplanations for Time Series forecasting (PAX-TS), a technique for multi-granular explanations, specific to time series forecasting. PAX-TS uses locally perturbed inputs to assess how a forecasting model generates its forecasts, giving insights at different levels of granularity, up to the time-step-level. Fig. 1 shows an example of input perturbations ($\mathbf{x}^{\prime}$) of the global minimum of an individual sample ($\mathbf{x}$) and the respective forecasts ($\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime}$) for each perturbed input with different scale parameters (+5% and -10%).
@@
    Method: We introduce PAX-TS, a post-hoc model-agnostic XAI technique designed for time series forecasting. PAX-TS can be used on both univariate and multivariate time series, and its explanations can be generated at different granularities, ranging from high-level cross-channel correlations to low-level time-step-importance explanations.
@@
    Evaluation: We benchmark 7 forecasting models on 10 datasets and compare PAX-TS to TS-MULE and ShapTime, then distill six visually distinct temporal-dependency patterns that align with model performance.
@@
    Practicality: We analyze runtime, complexity, and fidelity, showing that PAX-TS is suitable for real-time use while accurately characterizing the underlying forecasting model. Code is available at https://anonymous.4open.science/r/pax-ts-6410.
@@ II Background
Explainability for time series forecasts can be achieved in two ways: Inherently explainable models are algorithms that are designed to be more comprehensible, providing both the forecast and additional information, which explains the forecast in different ways. Secondly, post-hoc methods can be applied based on the trained model and its forecast to explain how the model has generated the respective forecast. baur2024explainability provide a meta-review on explainability for electric load forecasting, summarising a range of techniques used in this application area. In this section, we describe both inherently explainable forecasting algorithms and post-hoc explanation techniques to give a broad overview of the related work.
@@
Algorithms providing inherent explanations. N-BEATS  [oreshkin2019n] is a state-of-the-art univariate time series forecasting algorithm based on multilayer perceptron stacks, which additionally offers a more interpretable architecture. In this architecture, the provided explanations are based on trend and seasonality, which sum up to the model’s forecast. This simple yet effective approach allows the end-user to get a more fine-grained understanding of the model’s forecast by dividing the forecast into components, which are simpler to comprehend. lim2021temporal introduce temporal fusion transformers, an attention-based forecasting algorithm utilizing past data, static covariates, and known future inputs to make multivariate forecasts. Their algorithm provides values of relative importance to each of the input features. This results in an overall comparison, demonstrating which feature affects the forecast the most. However, this only explains which feature affects a forecast, whether in a positive or negative way, only in a specific segment, or for certain peaks. In a similar way, pantiskas2020interpretable use the attention mechanism to generate feature importance scores by time step for their proposed forecasting model. With this method, a specific output time step can be related to the entire input window, showing the importance of each input time step for this specific time step. This forecast explanation method is cumbersome, especially with multivariate data and longer input and output window lengths.
@@
Post-hoc explanation techniques. zhang2023shaptime propose ShapTime, a method that adapts Shapley Values [shapley1953value] to characterize the importance of segments in the input sequence. With this method, the importance of each input segment on the mean of the forecast can be assessed, effectively describing how much a given segment affects the forecast overall. While this is useful for some use cases, the granularity of the explanation is too coarse to provide actionable insights. troncoso2023new propose a method for discovering association rules for univariate time series forecasting. The resulting association rules link single values in the input and the output window with intervals of varying size. This type of explanation is problematic, as many association rules are found, and, in addition to that, their intervals are often large, providing little practical value for the end-user. TS-MULE [schlegel2021ts] is a post-hoc explanation technique for the time series forecasting task, utilizing the perturbation-based approach of LIME [ribeiro2016should] to make explanations. TS-MULE includes different segmentation techniques to generate explanations, resulting in feature importance scores for segments of the input window. This approach faces problems similar to other feature importance-based techniques like SHAP or attention maps, as the importance of a feature relates to the accuracy of the entire forecast, which is not sufficient to explain time series forecasting models.
@@ III The PAX-TS algorithm
For the multivariate time series forecasting problem, a $d$-dimensional time series subsequence of length $b$, $\mathbf{x}\in\mathbb{R}^{d\times b}$, is to be forecast for a horizon of length $h$, resulting in the forecast $\hat{\mathbf{y}}\in\mathbb{R}^{d\times h}$, aiming to approximate the ground truth $\mathbf{y}\in\mathbb{R}^{d\times h}$. Note that $\mathbf{x}$ and $\mathbf{y}$ are consecutive, and the problem is referred to as univariate time series forecasting when $d=1$. In the literature, $b$ and $h$ vary, where common scenarios range from short horizons $h\leq 48$ to longer horizons of up to $h=512$. Given an input subsequence $\mathbf{x}$, a trained forecasting model $M_{\theta}(\cdot)$ generates a forecast $\hat{\mathbf{y}}$, with more accurate forecasting models yielding a closer approximation to the ground truth $\mathbf{y}$, resulting in lower forecasting error. Our proposed approach to explaining time series forecasts, PAX-TS, is perturbation-based and focuses on the relation between input and output subsequences on an index level based on summary statistics and indices of interest. Table I gives an overview of all symbols used in this paper.
Symbol Meaning
$\mathbf{x}$ Input subsequence
$b$ Input length
$d$ Dimensionality
$\hat{\mathbf{y}}$ Forecast
$\mathbf{y}$ Ground Truth
$h$ Forecast horizon
$M_{\theta}(\cdot)$ Forecasting model
$\alpha$ Scale parameter
$\mathcal{A}$ Set of scale parameters
$\phi$ Perturbation parameters
$w$ Perturbation window size
$s$ Gaussian softness
$w_{p},w_{a}$ Positional, amplitude weight
$\mu$ Mean
$k$ Seasonal period length
$m$ Linear polynomial slope
$a$ Linear polynomial intercept
$\mathbf{x}^{\prime},\hat{\mathbf{y}}^{\prime}$ Perturbed input and forecast
$\Delta$ Distance between $\mathbf{x}$ and $\mathbf{x}^{\prime}$
$\pi(\cdot)$ Property of interest function
$\bar{r}_{\pi}$ Mean change ratio for property
$\bar{R}$ Cross-channel correlations
TABLE I: List of symbols
@@ III-A Localized perturbation by index
A localized Gaussian-smoothed perturbation is applied to perturb a positive time series subsequence $\mathbf{x}$ of length $b$ at a given time step $t$, denoted as $x_{t}$, by a scale parameter $\alpha$, where $\alpha>0$ indicates a shift in upward direction, and $\alpha<0$ indicates a shift in downward direction. We introduce this smoothing operation to achieve more plausible perturbations compared to point-based perturbations. Additionally, a width parameter $w\in\mathbb{N}$ affects the perturbation’s range, and a softness parameter $s\in\mathbb{R}$ influences the Gaussian fall-off. To apply the perturbation, we first calculate a positional weight $w_{p}$ for Gaussian-like smoothing, calculated for each time step $i$ of the subsequence. The weight is based on the distance between time step $i$ and $t$, and it is set to 0 for all time steps outside the fixed window size $w$:
@@
$w_{a}\;=\;\frac{\min(x_{i},x_{t})}{\max(x_{i},x_{t})+\epsilon},\qquad\forall\,i\in\{1,\dots,b\}$
@@
$\mathbf{x}^{\prime}\;=\;\{\,x_{i}^{\prime}\,\}_{i=1}^{b},\qquad x_{i}^{\prime}\;=\;x_{i}\;+\;w_{p}\,w_{a}\,\alpha\,x_{i},\quad\forall i\in\{1,\dots,b\}$
@@ III-B Summary statistic scaling
In this work, we consider the first and second moments of the time series as relevant summary statistics for scaling due to their simplicity and practical relevance. The same procedure can be followed for higher-order moments, however, the equations need to be adapted accordingly. With a scale parameter $\alpha$, the first moment of a time series subsequence $\mathbf{x}$ can be scaled as follows:
$\mathbf{x^{\prime}\;=\;\mathbf{x}+\alpha\mu},\qquad\mu=\frac{1}{b}\sum_{i=1}^{b}x_{i}$
@@
$\mathbf{x}^{\prime}\;=\;(\mathbf{x}-\mu)*\sqrt{\alpha+1}+\mu,\qquad\mu=\frac{1}{b}\sum_{i=1}^{b}x_{i}$
@@ III-C Trend-drift adjustment
The drift, representing the overall trend of a time series subsequence $\mathbf{x}$, can be adjusted by deseasonalizing the sequence, fitting a linear polynomial, and adjusting its slope according to a scale parameter $\alpha$. Given a seasonality length $k\in\mathbb{N}$, a discrete convolution operation, denoted by ”*”, results in a deseasonalized moving average of the time series:
@@
Using ordinary least squares, a linear polynomial can be fit to the residual $\mathbf{x}_{d}$, resulting in an intercept $a$ and a slope $m$, representing the trend of the deseasonalized sequence. The adjusted slope can be obtained as follows, while not adjusting the intercept, which provided the highest performance during early experiments: